# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v1

@@ Abstract
While 3D instance segmentation has made significant progress, current methods struggle to address realistic scenarios where new categories emerge over time with natural class imbalance. This limitation stems from existing datasets, which typically feature few well-balanced classes. Although few datasets include unbalanced class annotations, they lack the diverse incremental scenarios necessary for evaluating methods under incremental settings. Addressing these challenges requires frameworks that handle both incremental learning and class imbalance. However, existing methods for 3D incremental segmentation rely heavily on large exemplar replay, focusing only on incremental learning while neglecting class imbalance. Moreover, frequency-based tuning for balanced learning is impractical in these setups due to the lack of prior class statistics. To overcome these limitations, we propose a framework to tackle both Continual Learning and class Imbalance for 3D instance segmentation (CLIMB-3D). Our proposed approach combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module. Unlike prior methods, our framework minimizes ER usage, with KD preventing forgetting and supporting the IC module in compiling past class statistics to balance learning of rare classes during incremental updates. To evaluate our framework, we design three incremental scenarios based on class frequency, semantic similarity, and random grouping that aim to mirror real-world dynamics in 3D environments. Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline. Code is available at: https://github.com/vgthengane/CLIMB3D
@@ 1 Introduction
Figure 1: Class-incremental setup for 3D instance segmentation. As tasks progress through time, new classes are introduced incrementally. After each new task, the model should recognize both previously learned and newly introduced classes. For example, at Task-2, new classes such as Pillow, Coffee Table, and Sofa Chair are added, and the model is able to detect these three classes along with previous ones like Table, Chair, and Couch.
@@
3D instance segmentation is an essential task in computer vision that involves identifying and segmenting individual objects in the real physical space, playing a key role in applications across graphics, robotics, and autonomous systems. Its ability to provide precise object boundaries and class labels enhances scene understanding, facilitates object manipulation, and improves perception in dynamic environments.
@@
In recent years, a variety of methods have been proposed, including top-down approaches [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up approaches [hou20193dsis, yang2019learning], and transformer-based architectures [Schult23ICRA]. These methods have shown impressive results in traditional setups, which assume that all object classes are available during training. However, this assumption limits its applicability in real-world scenarios where new categories gradually emerge over time, often exhibiting naturally imbalanced distributions. Thus, there is a need for class-incremental learning (CIL) frameworks that not only adapt to new classes but also preserve prior knowledge, especially for rare or less frequent categories, which are more prone to catastrophic forgetting.
@@
Most existing research in class-incremental learning focuses on 2D image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with some extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. These methods employ strategies such as exemplar replay [buzzega2020dark, rebuffi2017icarl, cha2021co2l, kamra2017deep], regularization [aljundi2018memory, li2017learning, serra2018overcoming], and knowledge distillation [douillard2020podnet, kang2022class] to preserve previously learned knowledge and mitigate catastrophic forgetting [mccloskey1989catastrophic]. Few studies have applied CIL to point clouds; however, they mostly focus on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. At the scene level, some works have explored 3D semantic segmentation [Yang_2023_CVPR] with incremental learning, but their performance is not as competitive as state-of-the-art methods that do not employ incremental learning, which limits their applicability. Other methods tackle open-world incremental learning [boudjoghra20243d] but rely heavily on large exemplar sets [rolnick2019experience] and often neglect the challenge of class imbalance.
To address this, we propose CLIMB-3D, a unified framework that combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module to tackle Continual Learning for Imbalance 3D instance segmentation in indoor environments, as shown in Figure 1. Our framework operates as follows: ER stores a subset of representative samples from previous stages, allowing for effective replay during new task learning. KD transfers knowledge by retaining a copy of the model from the previous task, thereby mitigating forgetting. The IC module is specifically designed to reduce forgetting of rare classes by leveraging the frequency of object occurrence. However, since we do not have access to the previous task data or statistics during incremental phases, we instead compile these statistics from the previous model used by KD to generate weight for earlier categories. These weights are used to favor both frequent and rare classes, ensuring a balanced learning and mitigating forgetting.
@@
To evaluate CLIMB-3D in a realistic, incremental learning setup, we create three benchmark scenarios using the ScanNet200 dataset [rozenberszki2022language], which features 200 classes with natural class imbalances. These scenarios are designed to reflect real-world conditions where new categories emerge gradually and follow inherent class imbalances. These are based on 1 frequency of object occurrence, 2 semantic similarity between the object, and 3 random grouping. Our experiments demonstrate that CLIMB-3D significantly improves performance by effectively mitigating the forgetting of previous tasks compared to earlier exploration in class incremental 3D segmentation.
In summary, our contributions are as follows:
  1. 1.
    We propose a new problem setting for imbalanced class incremental 3D segmentation, along with a simple yet effective method to address this challenge by minimizing catastrophic forgetting and balancing the learning process.
  2. 2.
    To benchmark this setting, we design three scenarios aimed at simulating real-world conditions where objects emerge continuously with natural class imbalance.
  3. 3.
    Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline.
@@ 2.2 Incremental Learning
3 Preliminaries
3.1 Transformer-based Segmentation
We adopt a transformer-based instance segmentation method based on Mask3D [Schult23ICRA]. Specifically, transformer-based segmentation model $\Phi$ is employed for point cloud instance segmentation. Given an input point cloud $p$, the model predicts $\hat{y}=\{(\hat{m_{j}},\hat{c_{j}})\}_{j=1}^{J}$, which consists of mask predictions and class probabilities for each instance. The segmentation process begins by quantizing the input point cloud $p$ into voxels $V$, creating voxelized representations of size $\mathbb{R}^{M_{0}\times 3}$. Each voxel is assigned an average RGB color computed from the points within that voxel, serving as its initial feature representation. The feature backbone network generates a high-resolution output feature map $\mathbf{F}_{0}\in\mathbb{R}^{M_{0}}$. Additionally, intermediate feature maps are extracted from the decoder layers of the backbone network. For each intermediate feature map ($r\geq 0$), a set of $K_{r}$ voxels is selected, and their features are linearly projected to a fixed dimension $D$, yielding feature maps $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$.
The Transformer decoder initiates with a set of K instance queries and iteratively improves them using L Transformer decoder layers. These layers employ cross-attention to refine the instance queries, incorporating information from point cloud features. The decoder attends to a specific feature map obtained from the corresponding feature backbone layer at each layer, employing conventional cross-attention mechanisms. This process enables the decoder to reason at the instance level through self-attention, resulting in the generation of accurate and contextually relevant instance queries tailored to the specific scene.
To achieve this, the voxel features $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$ are transformed into sets of keys $\mathbf{K}\in\mathbb{R}^{M_{r}\times D}$ and values $\mathbf{V}\in\mathbb{R}^{M_{r}\times D}$ through linear projection. The instance queries $\mathbf{Z}$ are also projected to create the queries $\mathbf{Q}$. This enables cross-attention, allowing the queries to gather relevant information from the voxel features. Following cross-attention, a self-attention step occurs among the queries, facilitating information exchange and refinement. The learned queries are then used to make $K$ class and mask predictions, which are matched with ground truth labels through bipartite matching, resulting in $\hat{y}=\{(\hat{m}_{j},\hat{c}_{j})\}_{j=1}^{J}$. The model is optimized based on the ground truth label, mask, and class predictions:
$\mathcal{L}_{\mathrm{Seg}}(y_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m_{j},\hat{m}_{j})+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(c_{j},\hat{c}_{j})$ (1)
where, mask loss $\mathcal{L}_{\mathrm{mask}}=\lambda_{\mathrm{ce}}\mathcal{L}_{\mathrm{ce}}(y_{j},\hat{y}_{j})+\lambda_{\mathrm{dice}}\mathcal{L}_{\mathrm{dice}}(y_{j},\hat{y}_{j})$ and $\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}$ is classification loss.
The traditional setup assumes all categories are available and well-balanced during training. However, in scenarios where only a subset of categories is present, training the model in multiple phases is required. Unfortunately, such multi-phase models often suffer from forgetting previous tasks. To address this issue, we employ incremental learning strategies which will be discussed in the next section.
@@ 4 Methodology
In this section, we begin by formulating the incremental 3D instance segmentation problem and introduce our proposed method to address it using a transformer-based model. We then detail the design of incremental scenarios, outlining the motivation and considerations behind each one.
Figure 2: CLIMB-3D: At phase $t$ (for $t>1$), we create a copy of the model from the previous phase, $\Phi^{\mathrm{T}}$ (the teacher model), and designate it as $\Phi^{\mathrm{S}}$ (the student model). Both $\Phi^{\mathrm{T}}$ and $\Phi^{\mathrm{S}}$ process the input point cloud simultaneously, producing predictions $\hat{y}_{\mathrm{pseudo}}$ and $\hat{y}$, respectively. To prevent the model from forgetting less frequent categories from previous tasks, we balance the predictions from $\Phi^{\mathrm{T}}$, resulting in a balanced output $\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}$. The concatenated vector of ground truth labels $[y,\hat{y}_{\mathrm{pseudo}},\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}]$ is then compared with the predicted labels $\hat{y}$. A loss function is applied based on this comparison, enabling the student model to learn from the differences between the predicted and the concatenated pseudo-ground truth labels

### v2

@@ Abstract
While 3D instance segmentation (3DIS) has advanced significantly, existing methods typically assume that all object classes are known in advance and are uniformly distributed. However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance. Although some approaches have addressed class emergence, they often overlook class imbalance, resulting in suboptimal performance – particularly on rare categories. To tackle this challenge, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS. Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under constrained memory conditions. To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen prior model. Despite its promise, PLG tends to bias towards frequent classes. Therefore, we propose a class-balanced re-weighting (CBR) scheme, that estimates object frequencies from pseudo-labels and dynamically adjusts training bias – without requiring access to past data. We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset, and additionally on semantic segmentation on ScanNetV2. Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalization across both frequent and rare classes.
@@ 1 Introduction
Figure 1: Overview of the CI-3DIS setting: new classes are introduced incrementally as tasks progress. After each task, the model must recognize both previously learned and newly added classes. For instance, at Task 2, classes like Pillow, Coffee Table, and Sofa Chair are introduced, while the model retains the ability to detect earlier classes such as Table, Chair, and Couch.
@@
3D instance segmentation (3DIS) is an essential task in computer vision with widespread applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified]. It involves identifying and segmenting individual objects in 3D space, enabling precise object boundaries and class labels that enhance scene understanding and perception in dynamic environments.
@@
In recent years, a variety of methods for 3DIS have been proposed, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA]. These methods have demonstrated strong performance under the traditional setting, which assumes that all object classes are available and well-balanced during training. However, these assumptions limit their applicability in real world settings, where new categories gradually emerge over time, often exhibiting natural class imbalance.
@@
These limitations in current 3DIS methods underscore the need for a Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual] framework capable of learning new categories as they appear, while maintaining accuracy on previously learned classes – especially rare ones, which are more susceptible to without catastrophic forgetting [mccloskey1989catastrophic]. Unfortunately, most existing research in CIL has focused on the 2D domain for image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with limited extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. Extensions to 3D point clouds remain limited, with most work focusing on object-level classification rather than scene-level segmentation [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but typically rely on large exemplar buffers [rolnick2019experience] Moreover, they typically overlook class imbalance, limiting real-world applicability.
To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance. The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay. However, this alone does not yield promising results under the strict memory constraints needed CI-3DIS. To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen previous task model to generate supervision for earlier classes. However, we observed that, PLG tends to favour frequent classes while ignoring under-represented ones. To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data. Together, these components form a practical and effective solution for real-world 3DIS.
@@
To evaluate CLIMB-3D in a realistic incremental setup, we design three benchmark scenarios using the ScanNet200 [rozenberszki2022language] dataset, which contains 200 classes with a natural long-tail distribution. These scenarios simulate real-world conditions where new classes gradually emerge, following inherently imbalanced distributions. Specifically, new classes appear in groups based on one of three criteria: (A) object frequency, (B) semantic similarity, or (C) random grouping without a specific pattern. Our experiments on these scenarios for incremental 3DIS show that CLIMB-3D effectively mitigates forgetting of previous tasks while improving performance across both frequent and rare classes, delivering consistent gains over prior methods.
In summary, our contributions are threefold: (i) We introduce a novel problem setting of imbalanced class-incremental 3DIS, along with a simple and effective method to mitigate catastrophic forgetting and ensure balanced learning; (ii) We establish three benchmark scenarios that realistically model continuous object emergence with natural class imbalance; and (iii) We validate our framework’s effectiveness, achieving SOTA results with improvements up to 16.76% mAP over existing baselines [^1] Code will be available at: https://github.com/vgthengane/CLIMB3D .
@@ 3 Methodology
Figure 2: Overview of CLIMB-3D for CI-3DIS. The model incrementally learns new classes in sequential phases using ER to mitigate forgetting, a PLG for supervision on old classes, and CBR to handle class imbalance.